THe expressions 3x+1 and x+3 represent the length of two sides of a regular hexagon. FInd the value of x.
Answer by cleomenius(959)

If they are sides of a regular hexagon, the sides are equal.
3x + 1 = x + 3
2x = 2
x = 1
3(1) + 1 = 1 + 3.
4 = 4
This does check.
Cleomenius.